\begin{thebibliography}{10}

\bibitem{borger11}
James Borger.
\newblock The basic geometry of {W}itt vectors, {I}: {T}he affine case.
\newblock {\em Algebra $\&$ Number Theory}, 5(2):231--285, 2011.

\bibitem{bds1}
James Borger and Bart de~Smit.
\newblock Lambda actions of rings of integers.
\newblock {\em arXiv:1105.4662}, 2011.

\bibitem{bds2}
James Borger and Bart de~Smit.
\newblock Explicit class field theory and the algebraic geometry of
  $\lambda$-rings.
\newblock {\em arXiv:1809.02295}, 2018.

\bibitem{bc}
Jean-Benoit Bost and Alain Connes.
\newblock Hecke algebras, type {III} factors and phase transitions with
  spontaneous symmetry breaking in number theory.
\newblock {\em Selecta Mathematica}, 1(3):411--457, 1995.

\bibitem{cc14}
Alain Connes and Caterina Consani.
\newblock On the arithmetic of the {BC}-system.
\newblock {\em Journal of Noncommutative Geometry}, 8(3):873--945, 2014.

\bibitem{cc21}
Alain Connes and Caterina Consani.
\newblock Segal’s {G}amma rings and universal arithmetic.
\newblock {\em The Quarterly Journal of Mathematics}, 72(1-2):1--29, 2021.

\bibitem{cc23}
Alain Connes and Caterina Consani.
\newblock {BC}-system, absolute cyclotomy and the quantized calculus.
\newblock {\em EMS Surveys in Mathematical Sciences}, 9(2):447--475, 2023.

\bibitem{ccmot}
Alain Connes, Caterina Consani, and Matilde Marcolli.
\newblock Noncommutative geometry and motives: the thermodynamics of
  endomotives.
\newblock {\em Advances in Mathematics}, 214(2):761--831, 2007.

\bibitem{ccm}
Alain Connes, Caterina Consani, and Matilde Marcolli.
\newblock Fun with {$\mathbb F_1$}.
\newblock {\em Journal of {N}umber {T}heory}, 129(6):1532--1561, 2009.

\bibitem{cmr}
Alain Connes, Matilde Marcolli, and Niranjan Ramachandran.
\newblock {KMS} states and complex multiplication.
\newblock {\em Selecta Mathematica}, 11(3):325, 2006.

\bibitem{hp}
Eugene Ha and Fr{\'e}d{\'e}ric Paugam.
\newblock {B}ost-{C}onnes-{M}arcolli systems for {S}himura varieties. {P}art
  {I}. {D}efinitions and formal analytic properties.
\newblock {\em International Mathematics Research Papers}, 2005(5):237--286,
  2005.

\bibitem{lln}
Marcelo Laca, Nadia~S Larsen, and Sergey Neshveyev.
\newblock On {B}ost-{C}onnes type systems for number fields.
\newblock {\em Journal of {N}umber {T}heory}, 129(2):325--338, 2009.

\bibitem{yalkinoglu}
Bora Yalkinoglu.
\newblock On arithmetic models and functoriality of {B}ost-{C}onnes systems.
\newblock {\em Inventiones mathematicae}, 191(2):383--425, 2013.

\end{thebibliography}

\documentclass[11pt, a4paper]{amsart}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage[figuresright]{rotating}
\usepackage[all]{xy}
\usepackage{pdfsync}
\usepackage[toc,page]{appendix}
\usepackage{hyperref}
\usepackage{mathtools}
\usepackage{lipsum}
\usepackage[T1]{fontenc}        
\usepackage[utf8]{inputenc}
\usepackage[toc,page]{appendix}
\usepackage{tikz-cd}
\usepackage{accents}
\usepackage{fullpage}

\newcommand{\complex}{\mathbb{C}}

\newcommand{\okhat}{\widehat{\mathcal O }_K}
\newcommand{\ok}{{{\mathcal O }_K}}
\newcommand{\drf}{{\mathrm{DR}_\mathfrak f}}

\newcommand{\kab}{\mathrm{Gal}(K^{\mathrm{ab}}/K)}

\newcommand{\eq}[1]{\begin{equation}#1 \end{equation}}

\newcommand{\eqst}[1]{\begin{equation*}#1\end{equation*}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem*{theorem*}{Theorem}
\newtheorem{proposition}{Proposition}[section]

\newtheorem{corollary}{Corollary}[section]

\newtheorem{remark}{Remark}[section]

\numberwithin{equation}{section}

\begin{document}
\author{Bora Yalkinoglu}
\address{CNRS and IRMA, Strasbourg}
\email{yalkinoglu@math.unistra.fr}
\date{\today}
\title{Bost-Connes systems and periodic Witt vectors}
\maketitle

\begin{abstract}
\noindent
Using Borger's theory of periodic Witt vectors, we construct integral refinements of the arithmetic subalgebras associated with Bost-Connes systems for general number fields.

\end{abstract}

\section{Introduction}
\noindent Bost-Connes systems were introduced in the seminal paper \cite{bc} and have since been studied extensively; see, e.g., \cite{hp,cmr,lln,ccm,yalkinoglu,cc14}. Today, for each number field $K$, we can associate a Bost-Connes system $\mathcal A _K=(A_K,\sigma_t)$, a $C^{\mathrm{*}}$-dynamical system with the following key properties (among others):
\begin{itemize}
\item[(i)] The partition function of $\mathcal A_K$ is given by the Dedekind zeta function of $K$. \smallskip
\item[(ii)] The maximal abelian Galois group $\kab$ of $K$ acts by symmetries on $\mathcal A_K$. \smallskip
\item[(iii)] There exists a $K$-rational subalgebra $A_K^{\mathrm{arith}} \subset A_K$ such that, for every extremal $\mathrm{KMS}_\infty$-state $\varrho$ and every $f \in A_K^{\mathrm{arith}}$, the values $\varrho(f) \in K^{\mathrm{ab}}$ generate $K^{\mathrm{ab}}$ over $K$.
\item[(iv)] For $\nu \in \kab$ and $f \in A_K^{\mathrm{arith}}$, the following compatibility relation holds: \eqst{{}^{\nu}\hspace{-0.6mm}\varrho(f)  = \nu^{-1}(\varrho(f))\,.} 
\item[(v)] The $\complex$-algebra $A_K^{\mathrm{arith}}$$\otimes_K$$\complex$ is dense in $A_K$. \bigskip
\end{itemize}
The arithmetic subalgebras $A_K^{\mathrm{arith}}$ were constructed in full generality in \cite{yalkinoglu}, based on a Grothendieck-Galois correspondence for $\Lambda_K$-rings developed in \cite{bds1}. In the language of \cite{ccmot}, they give rise to algebraic endomotives $\mathcal E _K = A^{\mathrm{arith}}_K \rtimes I_K$, which serve as algebraic incarnations of the Bost-Connes systems $\mathcal A _K$. In this note, we use Borger's beautiful theory of periodic Witt vectors $\mathbb W_K^{\mathrm{(\mathfrak f)}}$, developed in \cite{bds2}, to construct an integral model\footnote{This was promised in \cite{yalkinoglu}.} of the arithmetic subalgebra $A_K^{\mathrm{arith}}$. Specifically, we construct an $\mathcal O _K$-algebra $A_K^{\mathrm{int}} \subset A_K^{\mathrm{arith}}$, satisfying analogous properties, such that $A_K^{\mathrm{int}} \otimes_{\mathcal O_K} K \cong A_K^{\mathrm{arith}}.$
\begin{theorem}
\label{integralmodel}
For every number field $K$, the $\mathcal O _K$-algebra $A_K^{\mathrm{int}} = \varinjlim_{^{\mathfrak f}} \mathbb W_K^{(\mathfrak f)}$ provides an integral model for the arithmetic subalgebra $A_K^{\mathrm{arith}}$ and gives rise to an integral refinement $\mathcal E^{\mathrm{int}}_K = A_K^{\mathrm{int}}\rtimes I_K$ of the algebraic endomotive $\mathcal E _K $ constructed in  \cite{yalkinoglu}.
\end{theorem} 
\noindent 
The ongoing interest in the interplay between BC-systems and Witt vectors (cf. \cite{cc14,cc21,cc23}) motivates the present note. Until now, it was known only that the original Bost–Connes system $\mathcal{A}_\mathbb{Q}$ could be expressed in terms of Witt vectors (see \cite{cc14}). Our main result demonstrates that, in fact, all Bost–Connes systems $\mathcal A _K$ admit a natural description in terms of (generalized) Witt vectors. Below, we outline the key ingredients involved in constructing our integral refinement $\mathcal{E}_K^{\mathrm{int}}$.

\section{The arithmetic subalgebra}
\noindent Let $K / \mathbb Q$ be a number field. Denote its ring of integers by $\mathcal O _K$, the monoid of (non-zero) integral ideals by $I_K$ and the subset of totally positive elements by $K_+$. Let $\okhat$ denote the profinite completion of $\ok$. For any ring $R$, write $R^\times$ for its group of units. \\
The goal of this section is to explain the construction of the arithmetic subalgebra $A_K^{\mathrm{arith}}$, using the Grothendieck-Galois correspondence for $\Lambda_K$-rings introduced by Borger and de Smit (see \cite{bds1}). We begin by recalling the necessary ingredients.
\subsection{$\Lambda_K$-rings}
We follow \cite{bds1,bds2}. For $\mathfrak p \in I_K$ a prime ideal, let $\kappa(\mathfrak p) = \mathcal O _K / \mathfrak p$ be the corresponding finite residue field. The Frobenius endomorphism $\mathrm{Frob}_\mathfrak p$ of a $\kappa(\mathfrak p )$-algebra is defined by $x \mapsto x^{\vert \kappa(\mathfrak p ) \vert}$. An endomorphism $\Psi$ of an $\mathcal O _K$-algebra $E$ is called a Frobenius lift at $\mathfrak p$ if $\Psi \otimes 1 = \mathrm{Frob}_\mathfrak p$ on $E \otimes_{\mathcal O _K} \kappa(\mathfrak p)$. \\ 
Let $E$ be an $\mathcal O _K$-algebra. A $\Lambda_K$-structure on $E$ is a family of endomorphisms $(\Psi_\mathfrak p)$, indexed by prime ideals $\mathfrak p \in I_K$, such that: \begin{itemize} \item[1)] $\Psi_\mathfrak p \circ \Psi_\mathfrak q = \Psi_\mathfrak q \circ \Psi_\mathfrak p$ for all $\mathfrak p,\mathfrak q$; \item[2)] each $\Psi_\mathfrak p$ is a Frobenius lift a $\mathfrak p$.
\end{itemize} 
We say that a $K$-algebra $E$ has an integral $\Lambda_K$-structure if there exists an $\mathcal O_K$-algebra $E'$ with $\Lambda_K$-structure such that $E \cong E' \otimes_{\mathcal O _K} K$. In this case, $E'$ is called an integral model of $E$. \\ 
A $\Lambda_K$-structure on an $\mathcal O_K$-algebra $E$ induces an action of $I_K$ on $E$ by $\mathcal O_K$-algebra endomorphisms; that is, we obtain a monoid map $I_K \to \mathrm{End}_{\mathcal O _K}(E)$.
\subsection{Deligne-Ribet monoid}
We follow \cite{bds1,bds2,yalkinoglu}. For $\mathfrak f \in I_K$, define the (finite) Deligne-Ribet monoid at $\mathfrak f$ as $\drf = I_K / \sim_{\mathfrak f}$, where the the equivalence relation is given by \eq{\mathfrak a \sim_\mathfrak f \mathfrak b :\Leftrightarrow \exists \, x \in K^\times_+ \cap (1+\mathfrak f \mathfrak b^{-1}) : (x) = \mathfrak a \mathfrak b^{-1}.}
More explicitly, one has the following descriptions: \eq{\label{key} \drf \overset{\text{\cite{bds1}}}{\cong} \coprod_{\mathfrak d \vert \mathfrak f} \mathrm{Gal}(K_{ \mathfrak d}/K) \overset{\text{\cite{yalkinoglu}}}{\cong} \mathcal O _K / \mathfrak f \times _{(\mathcal O _K / \mathfrak f)^\times} \mathrm{Gal}(K_\mathfrak f / K) ,} where $K \subset K_{\mathfrak d} \subset K^{\mathrm{ab}}$ denotes the strict ray class field of conductor $\mathfrak d$, and $K^{\mathrm{ab}}$ is the maximal abelian extension of $K$. For $\mathfrak f \, \vert \, \mathfrak f '$ there is a natural projection map $\pi_{\mathfrak f,\mathfrak f '} :\mathrm{DR}_{\mathfrak f'} \to \mathrm{DR}_{\mathfrak f}$, giving rise to the Deligne-Ribet monoid of $K$: \eq{\mathrm{DR}_K = \varprojlim_{\mathfrak f} \mathrm{DR}_\mathfrak f \overset{\text{\cite{yalkinoglu}}}{\cong} \okhat \times_{\okhat^\times} \mathrm{Gal}(K^{\mathrm{ab}}/K).}
\subsection{Grothendieck-Galois correspondence and the arithmetic subalgebra}
With these ingredients in place, we can now state the correspondence from \cite{bds1}:
\begin{theorem}[Grothendieck-Galois correspondence for $\Lambda_K$-rings]
\label{ggc}
The functor \eq{E \mapsto \mathrm{Hom}_{K-\mathrm{alg}}(E,\overline K)} induces a contravariant equivalence \eq{\mathfrak H_K : \mathcal E _{\Lambda,K} \longrightarrow \mathcal S _{\mathrm{DR}_K}} between the category $\mathcal E _{\Lambda,K}$ of finite, étale K-algebras with an integral $\Lambda_{K}$-structure and the category $\mathcal S _{\mathrm{DR}_K}$ of finite sets equipped with a continuous action of the Deligne-Ribet monoid $\mathrm{DR}_K$.
\end{theorem}
\noindent From (\ref{key}), we see that the finite, étale $K$-algebra $E_\mathfrak f = \mathfrak H_K^{-1}(\mathrm{DR}_\mathfrak f$) corresponds to the finite product of strict ray class fields: \eq{E_\mathfrak f = \prod_{\mathfrak d \, \vert \, \mathfrak f} K_{ \mathfrak d}.} 
The functoriality of $\mathfrak H_K$ gives rise to our arithmetic subalgebra \eq{A^{\mathrm{arith}}_K = \varinjlim_\mathfrak f E_\mathfrak f} and the compatibility of the $\Lambda_K$-structures on all the $E_\mathfrak f$ allows us to define the algebraic endomotive underlying the Bost-Connes system $\mathcal A _K$: \eq{\mathcal E _K =  A^{\mathrm{arith}}_K \rtimes I_K.}

\section{Periodic Witt vectors}
\noindent We follow Section 9 of \cite{bds2} and present the definitions relevant for constructing integral models our $E_\mathfrak f$. \\ 
The generalized Witt vectors $\mathbb W_K(R)$ of a (flat) $\ok$-algebra $R$ are defined as the maximal $\ok$-subring of the ghost ring $R^{I_K}$ that satisfies the following properties: \begin{itemize} \item[1)] It is stable under the natural action of $I_K$; \item[2)] each prime ideal $\mathfrak p \in I_K$ induces a Frobenius lift $\Psi_\mathfrak p :\mathbb W_K(R) \to \mathbb W_{K}(R)$ at $\mathfrak p$. 
\end{itemize}
\begin{remark}
This construction is functorial and recovers the classical notions of big Witt vectors and $p$-typical Witt vectors (see \cite{borger11}).
\end{remark}
\noindent For $\mathfrak f \in I_K$, the $\mathfrak f$-periodic Witt vectors of $R$ are defined by \eq{\mathbb W_K^{(\mathfrak f)} = \{ x \in \mathbb W_K(R) \, \vert \, \Psi_\mathfrak a (x) = \Psi_\mathfrak b(x) \  \forall \, \mathfrak a \sim_\mathfrak f \mathfrak b \}.} In particular, we can now consider the $\mathfrak f$-periodic Witt vectors of the integral closure $\overline{\mathcal O}_K$: \eq{\mathbb W_K^{(\mathfrak f)} = \mathbb W_K^{(\mathfrak f)}(\overline{\mathcal O}_K).}
The key observation for us is the following 
\begin{proposition}[\cite{bds2}]
There is an isomorphism of finite, étale $K$-algebras \eq{\mathbb W_K^{(\mathfrak f)} \otimes_\ok K \cong \prod_{\mathfrak d \vert \mathfrak f}K_{\mathfrak f / \mathfrak d}.}
\end{proposition}
\begin{corollary}[\cite{bds2}]
The periodic Witt vectors $\mathbb W_K^{(\mathfrak f)}$ provide an integral model for $E_\mathfrak f$, compatible with the respective $\Lambda_K$-structures.
\end{corollary}
\noindent The corollary immediately implies Theorem \ref{integralmodel}.

\bibliographystyle{plain}
\bibliography{bc_witt}

\end{document}